OpenStudy (anonymous):

find the mean of the discrete random variable with the probability distribution below x | 5 9 12 14 19 p(x)| 0.2 0.15 0.1 0.3 0.25

OpenStudy (anonymous):

multiply each x by each p(x) and add the result

OpenStudy (anonymous):

Mean is summation of [x times p(x)]

OpenStudy (anonymous):

\[5\times .2+9\times .15+12\times .1+14\times .3+19\times .25\]

OpenStudy (anonymous):

thats all i have to do @satellite

OpenStudy (anonymous):

that is all!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw